# Fast Transaction Scheduling in Blockchain Sharding

### v1

@@ 5.2. Analysis for Line Graph with Buckets
If the shards are in a line graph and each transaction accesses $k$ objects, then Algorithm 2 provides an $O(k\sqrt{d}\log D)$ approximation to the optimal schedule, where $d$ is the worst distance of transaction and its accessing objects and $D$ diameter of the line graph.
Proof.
From Lemma 8, time to process all the transactions in any bucket $B_{i}$ is at most $O(2^{i}kl_{i}\log D)$. Moreover, from Lemma 7, lower bound to process all the transaction of bucket $B_{i}$ is at least $2^{i}+l_{i}^{2}/4=O(2^{i}+l_{i}^{2})$, which gives us $O(\frac{2^{i}kl_{i}\log D}{2^{i}+l_{i}^{2}})$ approximation. we will analyze two cases:
(i) When $l_{i}<2^{i/2}$:
$O(\frac{2^{i}kl_{i}\log D}{2^{i}+l_{i}^{2}})\leq O(\frac{2^{i}k2^{i/2}\log D}{2^{i}+{2^{i/2}}^{2}})<O(\frac{2^{i}k2^{i/2}\log D}{2^{i}})=O(k\sqrt{2^{i}}\log D)\ .$
We can replace $2^{i}$ with $d$ because $2^{i}$ is the distance between the home shard of the transactions and its accessing objects, and $d$ is the maximum distance between transactions and their accessing objects in the line graph (L). Thus, we will get $O(k\sqrt{d}\log D)$.
(ii) When $l_{i}\geq 2^{i/2}$:
$O(\frac{2^{i}kl_{i}\log D}{2^{i}+l_{i}^{2}})\leq O(\frac{2^{i}kl_{i}\log D}{l_{i}^{2}})=O(\frac{2^{i}k\log D}{l_{i}})=O(k\sqrt{2^{i}}\log D)$
$=O(k\sqrt{d}\log D)\ .$
@@
From the above analysis and expression, we can say that the line graph Algorithm 1 gives an $O(k\sqrt{d}\log D)$ approximation to optimal schedule. ∎
6. Distributed Scheduler (DS)
@@
The scheduling algorithm we presented earlier uses a central authority in a shard with knowledge about all current transactions and the maximum degree of the transaction graph. However, such a central authority might not exist in blockchain sharding since each transaction is generated in a distributed manner in any shard. Here, we discuss a fully distributed scheduling approach using a clustering technique that allows the transaction schedule to be computed in a decentralized manner without requiring a central authority in the shard. We can also view distributed scheduling Algorithm 3 as a simulation of centralized scheduling Algorithm 1, which is described in Subsection 6.3.
6.1. Shard Clustering
In the distributed scheduler, shards are distributed across the network, and the distance between the home shard of the transaction and its accessing objects (destination shards) ranges from $1$ to $d\leq D$, where $D$ is the diameter of the shard graph. Let us suppose shards graph $G_{s}$ constructed with $s$ shards, where the weights of edges between shards denote the distances between them. We consider that $G_{s}$ is known to all the shards. We define $z$-neighborhood of shard $S_{i}$ as the set of shards within a distance of at most $z$ from $S_{i}$. Moreover, the 0-neighborhood of shard $S_{i}$ is the $S_{i}$ itself.
@@ 6.1. Shard Clustering
We consider that our distributed scheduling algorithm uses a hierarchical decomposition of $G_{s}$ which is known to all the shards and calculated before the algorithm starts. This shard clustering (graph decomposition) is based on the clustering techniques in (Gupta et al., 2006) and which were later used in (Sharma and Busch, 2014; Busch et al., 2022; Adhikari et al., 2024a, b). We divide the shard graph $G_{s}$ into the hierarchy of clusters with $H_{1}=\lceil\log D\rceil+1$ layers (logarithms are in base 2), and a layer is a set of clusters, and a cluster is a set of shards. Layer $q$, where $0\leq q<H_{1}$, is a sparse cover of $G_{s}$ such that:
  • Every cluster of layer $q$ has (strong) diameter of at most $O(2^{q}\log s)$.
@@
  • For each shard $S_{i}$ there exists a cluster at layer $q$ which contains $(2^{q}-1)$- neighborhood of $S_{i}$ within that cluster.
The clusters have a strong diameter, which measures distances within a cluster on the induced subgraph of $G_{s}$.
@@
For each layer $q$, the sparse cover construction in (Gupta et al., 2006) is actually obtained as a collection of $H_{2}=O(\log s)$ partitions of $G_{s}$. These $H_{2}$ partitions are ordered as sub-layers of layer $q$ labeled from 00 to $H_{2}-1$. A shard might participate in all $H_{2}$ sub-layers but potentially belongs to a different cluster at each sub-layer. At least one of these $H_{2}$ clusters at layer $q$ contains the whole $12^{q}-$ 1 neighborhood of $S_{i}$.
@@
In each cluster at layer $q$, a leader shard is specifically designated such that the leader’s $(2^{q}-1)$-neighborhood is in that cluster. As we give an idea of layers and sub-layers, we define the concept of height as a tuple $h=(h_{1},h_{2})$, where $h_{1}$ denotes the layer and $h_{2}$ denotes the sub-layer. Similar to (Sharma and Busch, 2014; Busch et al., 2022; Adhikari et al., 2024a, b), heights follow lexicographic order.
@@
The home cluster for each transaction $T_{i}$ is defined as follows: suppose $S_{i}$ is the home shard of $T_{i}$, and $z$ is the maximum distance from $S_{i}$ to the destination shards that will be accessed by $T_{i}$; the home cluster of $T_{i}$ is the lowest-layer (and lowest sub-layer) cluster in the hierarchy that contains $z$-neighborhood of $S_{i}$. Each home cluster consists of one dedicated leader shard, which will handle all the transactions that have their home shard in that cluster (i.e., transaction information will be sent from the home shard to the cluster leader shard to determine the schedule).
6.2. Distributed Scheduling Algorithm
@@
In the distributed scheduler Algorithm 3, each home shard determines the home cluster for its transaction based on the distance between the home shard of the transaction and its accessing objects (destination shards) and sends that transaction information to the leader of the respective home cluster. Each cluster $C_{z}$ belongs to some layer $q$ and sublayer $r$, where $0\leq q<H_{1}$ and $0\leq r<H_{2}$; Suppose a transaction $T_{i}$ with home cluster $C_{z}$; we also say that $T_{i}$ is at level $(q,r)$. The leader shard of $C_{z}$ assigns an integer color to each transaction $T_{i}$. We define the concept of the height for $T_{i}$ represented by a tuple $(q,r,\text{color})$. The heights of transactions are ordered lexicographically, and a the priority scheme is implemented. The priority for committing transactions is determined based on this order, giving precedence to transactions generated in lower cluster layers. The reason behind this priority scheme is to take the benefit of the locality to improve the transaction latency. Note that a home shard may have transactions at various levels. Those transactions will be processed at the respective level.
@@
For each level $(q,r)$, each cluster $C_{z}$ invokes the Algorithm 1 parallelly to schedule and process the transactions. While invoking Algorithm 1 for each cluster $C_{z}$ we set the parameters as $G[C_{z}]$, ${\mathcal{T}}(C_{z})$, $\ell(C_{z})$, where $G[C_{z}]$ is the cluster graph induced by $C_{z}$, ${\mathcal{T}}(C_{z})$ is the set of transactions whose home cluster is $C_{z}$, and $\ell(C_{z})$ is the leader shard of the cluster $C_{z}$ which is also presented in Algorithm 3.
@@
4Each cluster $C_{z}$ belongs at some level $(q,r)$, at layer $q$ and sublayer $r$ of the hierarchical decomposition, where $0\leq q<H_{1}$ and $0\leq r<H_{2}$; the levels $(q,r)$ are ordered lexicographically;
@@
8The home shard of each txn $T_{i}$ picks as a home cluster as the lowest level $(q,r)$ cluster $C_{z}$, which includes $T_{i}$ and the destination shards of its subtransactions;
@@
14       for each cluster $C_{z}$ in level $(i,j)$ in parallel do
15             ${\mathcal{T}}(C_{z})\leftarrow$ set of transactions with home cluster $C_{z}$;
16             $\ell(C_{z})$leftarrow leader shard of cluster $C_{z}$;
17             $G[C_{z}]\leftarrow$ cluster graph induced by $C_{z}$;
18             Run CentralizedScheduler($G[C_{z}]$, ${\mathcal{T}}(C_{z})$, $\ell(C_{z})$);
@@
6.3. Simulating Centralized Scheduler
@@
We can use the centralized scheduler algorithm (Algorithm 1) in each cluster of the distributed scheduler. In Algorithm 1, each leader shard has knowledge of transactions assigned to it and schedules them in a conflict-free manner using coloring. Similarly, in the distributed scheduler, each cluster leader shard knows all the transactions within its cluster and determines the transaction schedule by coloring the transactions conflict graph.
For example, let’s suppose the shard cluster graph has a maximum diameter $D$. Assume the distance between all the transactions and their accessed objects is $d\leq D$ far away from each other. In this scenario, each home shard of the transaction reports to the same home cluster, and the leader shard of that home cluster determines the schedule by coloring all the transactions similar to Algorithm 1. However, the distributed scheduler algorithm can provide significant benefits if transactions access nearby objects, allowing for parallel processing at each cluster level. Moreover, the distributed scheduler algorithm is based on hierarchical clustering, which needs to pay some constant factor delay due to layer and sublayer hierarchy with a priority scheme.
@@
Moreover, to improve the bounds for the Line graph and for special cases in the general graph when the $k$ objects are selected randomly in the distributed setting, we can simulate the Centralized Scheduler with Bucket (Algorithm 2) in a distributed manner. For each cluster $C_{z}$ graph (induced graph), we can run Algorithm 2, which will ultimately simulate Algorithm 1.
@@
Due to the layer and sublayer structure, we need to pay an $O(\log^{2}s)$ cost over the analysis obtained in the centralized approach, which is provided in the following Theorem 1.
Theorem 1.
@@
The distributed scheduler has a competitive ratio of $O(\mathcal{A}_{\mathcal{CS}}\log^{2}s)$, where $\mathcal{A}_{\mathcal{CS}}$ is the approximation ratio of the centralized scheduling algorithm ($\mathcal{CS}$).
Proof.
If the transactions are scheduled by the centralized scheduling algorithm 1, then their execution time would be within the centralized scheduler from optimal. However, in the distributed approach, the schedule is restricted by the layer and sublayer hierarchy of the cluster, as the lowest layer gets priority over the higher layer.
Suppose we only have transactions with cluster layer $q_{i}$ having cluster diameter $d_{q}$. Let $S_{q}$ be the leader shard of cluster $q_{i}$. Let ${\mathcal{T}}_{w}$ denote the set of transactions in $S_{q}$ that use object $O_{w}$. Let’s denote $l_{w}=|{\mathcal{T}}_{w}|$ and $l=\max_{w}l_{w}$ Suppose each transaction uses $k$ objects, and every object is used by at most $l$ transactions. Thus, the maximum degree in the transactions graph is at most $kl$, and therefore, the transactions graph can be greedily colored with $\xi\leq kl+1$ colors. Moreover, the diameter of the cluster is $d_{q}$, and for each color, Algorithm 1 takes at most $4$ rounds of interactions between the home shard of transaction and its accessing objects (destination shards). Thus, it takes $4d_{q}(kl+1)=O(kld_{q})$ rounds to process the transactions. Since $\Omega(l)$ is the lower bound, it gives $O(kd_{q})$ approximation. Which is equal to the approximation of the centralized scheduler algorithm Theorem 3. Thus we can say $\mathcal{A}_{\mathcal{CS}}=O(kd)$ for diameter $d$.
Let $q^{\prime}$ be the maximum layer accessed by any transaction where the diameter is at most $d_{q}^{\prime}$. If we only have layer $q_{i}$, the scheduler will require $4d_{q}(kl+1)$ rounds. However, if we consider the distributed scheduler with the layer and sublayer hierarchy, $S_{q}$ needs to process transactions from all layers $0,1,\cdots,q^{\prime}$ and sublayers indexed by $r$, which goes from 00 to $H_{2}-1$.
As cluster layer $q$ has diameter $O(2^{q}\log s)$, which is discussed in the shard clustering subsection above. So $d_{q}=O(2^{q}\log s)$ and we can write $d_{q}=c_{1}2^{q}\log s$ for some constant $c_{1}$. This implies $\sum_{q=0}^{q^{\prime}}d_{q}\leq 2d_{q}$. Thus, the total number of rounds needed to schedule all level $(q,r)$ clusters at $S_{q}$ is at most
(3) $\sum_{q=o}^{q^{\prime}}\sum_{r=0}^{H_{2}-1}4d_{q}(kl+1)\leq 8d_{q}(kl+1)H_{2}\ .$
We can replace $H_{2}=c_{1}^{\prime}\log s$ and $d_{q}=c_{2}^{\prime}d\log s$. Let’s combine positive constant $c_{1}^{\prime}$ and $c_{2}^{\prime}$ as $c$. The equation 3 becomes $8\cdot cd(kl+1)\log^{2}s$. Since $\Omega(l)$ is the lower bound and $\mathcal{A}_{\mathcal{CS}}=O(kd)$, we conclude that the distributed scheduler has competitive ratio $O(\mathcal{A}_{\mathcal{CS}}\log^{2}s)$. ∎
@@
From Theorem 1, we conclude that we need to pay an extra $O(\log^{2}s)$ rounds penalty for the distributed scheduler (Algorithm 3) due to hierarchical clustering. Hence, we have the following corollary:
@@
  • For a General graph, Algorithm 3 provides an $O(kd\log^{2}s)$ approximation to the optimal schedule, where $d$ is the maximum distance between the transaction and its accessing objects.
@@
  • Algorithm 3 provides an $O(k\log^{3}s)$ approximation for Hypercube, Butterfly, and $g$-dimensional grids graphs, where $g=O(\log s)$.
@@
  • For a General graph, If the transaction accesses $k$ objects randomly, then Algorithm 3 gives an $O(k\log D\cdot(k+\log s)\log^{2}s)$ approximation to the optimal schedule with probability at least $1-1/s$, where $D$ is the diameter of shard graph.

### v3

@@ V-B Analysis for Line Graph with Buckets
For the Line shard graph where each transaction accesses at most $k$ objects, Algorithm 2 provides an $O(k\sqrt{d}\log D)$ approximation to the optimal schedule.
Proof:
Let $B_{m}$, $0\leq m\leq\lfloor\log D\rfloor$, be the bucket that maximizes the term $2^{m}kl_{m}$. Since we have $\lfloor\log D\rfloor+1=O(\log D)$ buckets, from Lemma 8 the time to process the transactions in all the buckets is at most $O(2^{m}kl_{m}\log D)$. From Lemma 7, the time to process the transactions of bucket $B_{m}$ is at least $2^{m}+l_{m}^{2}/8=\Omega(2^{m}+l_{m}^{2})$. Therefore, we obtain an $O\left(\frac{2^{m}kl_{m}\log D}{2^{m}+l_{m}^{2}}\right)$ approximation factor for time. We analyze two cases:
(i) $l_{m}<2^{m/2}$:
$\frac{2^{m}kl_{m}\log D}{2^{m}+l_{m}^{2}}\leq\frac{2^{m}k2^{m/2}\log D}{2^{m}}=k\sqrt{2^{m}}\log D\ .$ (1)
By the definitions of $B_{m}$ and $d$, $2^{m+1}\leq d$. Thus, we get $O(k\sqrt{d}\log D)$ approximation factor.
(ii) $l_{m}\geq 2^{m/2}$:
$\frac{2^{m}kl_{m}\log D}{2^{m}+l_{m}^{2}}\leq\frac{2^{m}kl_{m}\log D}{l_{m}^{2}}=\frac{2^{m}k\log D}{l_{m}}\\ \leq k\sqrt{2^{m}}\log D=O(k\sqrt{d}\log D)\ .$ (2)
@@
Combining cases (i) and (ii) we get that Algorithm 2 gives an $O(k\sqrt{d}\log D)$ approximation to the optimal schedule. ∎
VI Distributed Scheduler (DS)
@@
The scheduling algorithms we presented earlier use a central authority in a shard that has knowledge about all current transactions and the maximum degree of the transaction graph. However, such a central authority might not exist in blockchain sharding since each transaction is generated in a distributed manner in any shard. Here, we discuss a fully distributed scheduling approach using a clustering technique that allows the transaction schedule to be computed in a decentralized manner without requiring a single central authority. Each cluster has its own leader and within each cluster the distributed approach invokes centralized scheduling Algorithm 1.
VI-A Shard Clustering
We define the $z$-neighborhood of shard $S_{i}$ as the set of shards within a distance of at most $z$ from $S_{i}$. Moreover, the 0-neighborhood of shard $S_{i}$ is just the $S_{i}$ itself.
@@ VI-A Shard Clustering
We consider that our distributed scheduling algorithm uses a hierarchical decomposition of $G_{s}$ which is known to all the shards and calculated before the algorithm starts. This shard clustering (graph decomposition) is based on the clustering techniques in [26] and which were later used in [43, 46, 16]. We divide the shard graph $G_{s}$ into the hierarchy of clusters with $H_{1}=\lceil\log D\rceil+1$ layers (logarithms are base 2). A layer is a set of clusters, where a cluster is a set of shards. Layer $q$, where $0\leq q<H_{1}$, is a sparse cover of $G_{s}$ such that:
  • Every cluster of layer $q$ has diameter of at most $O(2^{q}\log s)$.
@@
  • For each shard $S_{i}$ there exists a cluster at layer $q$ which contains the $(2^{q}-1)$-neighborhood of $S_{i}$ within that cluster.
@@
For each layer $q$, the sparse cover construction in [26] is actually obtained as a collection of $H_{2}=O(\log s)$ partitions of $G_{s}$. These $H_{2}$ partitions are ordered as sub-layers of layer $q$ labeled from 00 to $H_{2}-1$. A shard might participate in all $H_{2}$ sub-layers, and at least one of these $H_{2}$ clusters at layer $q$ contains the whole $12^{q}-$ 1 neighborhood of $S_{i}$.
@@
In each cluster at layer $q$, a leader shard is specifically designated such that the leader’s $(2^{q}-1)$-neighborhood is in that cluster. To express layers and sub-layers of a cluster, we define the concept of height as a tuple $h=(h_{1},h_{2})$, where $h_{1}$ denotes the layer and $h_{2}$ denotes the sublayer of the cluster. Similar to [43, 46, 16], heights are ordered lexicographically.
@@
The home cluster for each transaction $T_{i}$ is defined as follows: suppose $S_{i}$ is the home shard of $T_{i}$, and $z$ is the maximum distance from $S_{i}$ to the destination shards that will be accessed by $T_{i}$; the home cluster of $T_{i}$ is the lowest-layer (and lowest sub-layer) cluster in the hierarchy that contains the $z$-neighborhood of $S_{i}$. Each home cluster has a dedicated leader shard, which will handle all the transactions that have their home shard in that cluster (i.e., transaction information will be sent from the home shard to the cluster leader shard to determine the schedule).
VI-B Distributed Scheduling Algorithm
@@
In the distributed scheduler Algorithm 3, each home shard determines the home cluster for its transaction based on the distance between the home shard of the transaction and its accessing objects (destination shards) and sends that transaction information to the leader of the respective home cluster. Each cluster $C$ belongs to some layer $q$ and sublayer $r$ (height $(q,r)$), where $0\leq q<H_{1}$ and $0\leq r<H_{2}$. Suppose a transaction $T_{i}$ has home cluster $C$; we also say that $T_{i}$ is at height $(q,r)$. The leader shard of $C$, denoted $S_{\ell}(C)$, assigns an integer color to each transaction $T_{i}$. We define the concept of the height for $T_{i}$ represented by a tuple $(q,r,\text{color})$.
Using the lexicographic order of heights, we implement a priority scheme for the schedule. The priority for committing transactions is determined based on this order, giving precedence to transactions generated in lower cluster layers (and within a layer priority is given to lower sublayers). The reason behind this priority scheme is to take the benefit of the locality to improve the transaction latency. Note that a home shard may have transactions at various heights. Those transactions will be processed at their respective heights.
@@
For each level $(q,r)$, each cluster $C$ invokes Algorithm 1 parallelly to schedule and process the transactions. While invoking Algorithm 1 for each cluster $C$ we set the parameters as $G_{s}[C]$, ${\mathcal{T}}(C)$, $S_{\ell}(C)$, where $G_{s}[C]$ is the cluster graph induced by $C$ (i.e. containing the shards in $C$), ${\mathcal{T}}(C)$ is the set of transactions whose home cluster is $C$, and $S_{\ell}(C)$ is the leader shard of the cluster $C$.
@@
4Each cluster $C$ belongs at some height $(q,r)$, at layer $q$ and sublayer $r$ of the hierarchical decomposition, where $0\leq q<H_{1}$ and $0\leq r<H_{2}$; the heights $(q,r)$ are ordered lexicographically;
@@
8The home shard of each txn $T_{i}$ picks for home cluster the lowest height $(q,r)$ cluster $C^{\prime}$, which includes $T_{i}$ and the destination shards (which are at distance at most $z$) of its subtransactions;
@@
14       for each cluster $C$ in level $(q,r)$ in parallel do
15             ${\mathcal{T}}(C)\leftarrow$ set of transactions with home cluster $C$;
16             $S_{\ell}(C)$leftarrow leader shard of cluster $C$;
17             $G_{s}[C]\leftarrow$ cluster graph induced by $C$;
18             Run CentralizedScheduler($G_{s}[C]$, ${\mathcal{T}}(C)$, $S_{\ell}(C)$);
@@
VI-C Invocation of Centralized Scheduler
@@
The distributed scheduler (Algorithm 3) invokes the centralized scheduler (Algorithm 1) in each home cluster. By invoking Algorithm 1, the leader shard $S_{\ell}(C)$ in a cluster $C$ knows the transactions assigned to it and schedules them in a conflict-free manner using vertex coloring of $G_{s}[C]$.
@@
We can also use (Algorithm 2), instead of Algorithm 1, to further improve the bounds for the Line graph and for special cases in the general graph when the $k$ objects are selected randomly. Thus, for each cluster $C$, we can run Algorithm 2, which will ultimately invoke Algorithm 1.
@@
Next, we show that due to the layer and sublayer structure of the cluster hierarchy, we need to pay an $O(\log d\cdot\log s)$ factor cost over the approximation obtained in the centralized approach, which is proven in the following theorem.
Theorem 4
@@
The distributed scheduler (Algorithm 3) has a competitive ratio of $O(\mathcal{A}_{\mathcal{CS}}\cdot\log d\cdot\log s)$, where $\mathcal{A}_{\mathcal{CS}}$ is the approximation ratio of the centralized scheduling algorithm ($\mathcal{CS}$).
Proof:
Let $C$ be a cluster that takes the maximum time $\tau$ to process transactions compared to any other cluster in the hierarchy. Then, we can write
$\tau\leq\mathcal{A}_{\mathcal{CS}}\cdot\tau^{*}$
where $\tau^{*}$ is the optimal scheduling time for $C$.
Now, we consider all the $H_{1}$ layers and $H_{2}$ sublayers within each layer. The algorithm will only use transactions up to a layer $q^{\prime}=O(\log d)$, since the clusters at that layer cover all the transactions and the respective objects that they access. Thus, the total rounds required by Algorithm 3 to process all the transactions from all cluster level $(q,r)$ is at most
$\tau_{total}\leq\sum_{q=0}^{q^{\prime}}\sum_{r=0}^{H_{2}-1}\tau\leq\sum_{q=0}^{q^{\prime}}\sum_{r=0}^{H_{2}-1}\mathcal{A}_{\mathcal{CS}}\cdot\tau^{*}=\mathcal{A}_{\mathcal{CS}}\cdot\tau^{*}\cdot c\log d\cdot\log s$
for some constant $c>0$. Since $\tau^{*}$ is the lower bound for any scheduler, we can calculate the approximation ratio as:
$\frac{\tau_{total}}{\tau^{*}}\leq\frac{\mathcal{A}_{\mathcal{CS}}\cdot\tau^{*}\cdot c\log d\log s}{\tau^{*}}=O(\mathcal{A}_{\mathcal{CS}}\log d\cdot\log s)\ .$
@@
Since $\mathcal{A}_{\mathcal{CS}}$ is the approximation ratio of the centralized scheduling algorithm, from Theorem 4, we conclude that we need to pay an extra $O(\log d\cdot\log s)$ rounds penalty for the distributed scheduler (Algorithm 3) due to hierarchical clustering. Hence, we have the following results:
@@
  • For a General graph, Algorithm 3 provides an $O(kd\cdot\log d\cdot\log s)$ approximation to the optimal schedule.
@@
  • For Hypercube, Butterfly, and $g$-dimensional grids graphs, Algorithm 3 provides an $O(k\cdot\log d\cdot\log^{2}s)$ approximation to optimal schedule, where $g=O(\log s)$.
@@
  • For a General graph, If the transaction accesses $k$ objects randomly, then Algorithm 3 gives an $O(k\log D\cdot(k+\log s)\log d\cdot\log s)$ approximation to the optimal schedule with probability at least $1-1/s$.